OpenStudy (anonymous):

Randall has an account balance of $675.54 in a savings account that earns interest at a rate of 2.5% compounded twice a year. If Randall opened the account eleven years ago, what was the principal deposit rounded to the nearest dollar?

OpenStudy (anonymous):

\[A = P(1 + \frac{r}{n})^{nt}\]\[675.54 = P(1 + \frac{0.025}{2})^{2(11)}\]Just solve for P :)